# Perspectives on the arithmetic nature of the ratios $\zeta(2n+1)/\pi^{2n+1}$ and $\beta(2n)/\pi^{2n}$

### v1

@@ Introduction
Given the futility of the extraordinarily numerous attempts to compute these numbers, the solution of the designated problem is generally regarded as impossible; but so far a strong proof of this impossibility is lacking. Only heuristic methods and deep computations by brute force support this impossibility to this day. One may solve the problem by proving that the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ and $\dfrac{\beta(2n)}{\pi^{2n-1}}$ are not multiples of $\pi$. The difficulty of this new reformulation partly stems from the lack of sufficiently many closed-form expressions or identities involving these values that might serve as a starting point for an irrationality or transcendence proof. This paper is devoted to some integral representations of the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ and $\dfrac{\beta(2n)}{\pi^{2n}}$ through interesting logarithmic, trigonometric and hyperbolic kernels, along with several conjectures that support their non representability as rational multiples of $\pi$.
@@
The proof of this statement is fully detailed in [3] and is omitted here in the interest of succinctness. We merely note the existence of integrals $I_{n}$ satisfying
@@
In fact, the existence of integrals holding either the form (2.4) or (2.5) is the main and central result of this paper. The following table summarizes some of them, where $n\in\mathbb{N}^{*}$ and $C_{p,n}$ being rational coefficients with $p$ running over the set $\llbracket 1,n\rrbracket$
@@
We remark that the integral $\int_{0}^{1}\frac{\operatorname{Im}\!\left[\operatorname{Li}_{-2n}(ix)\right]}{x}\,\ln\ln\tfrac{1}{x}\,dx$ generalizes the integral
@@
that has been firstly evaluated by Adamchik [5] and rediscovered by Blagouchine [4]. Additionally, the integral $\int_{0}^{1}\frac{\operatorname{Li}_{-2n-1}(-x^{2})}{x}\,\ln\ln\tfrac{1}{x}\,dx$ exhibits a general form of $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ by also extending the Blagouchine’s integral $\int_{0}^{1}\dfrac{x(x^{4}-4x^{2}+1)\ln\ln\frac{1}{x}}{(1+x^{2})^{4}}dx=\frac{7\,\zeta(3)}{8\,\pi^{2}}$ to all higher $n$. The general formulae and closed-form expressions are provided by the following propositions.
@@
with rational coefficients. The following lines only give instructions on how to prove that
@@
holds form (2.4) and $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n+1}\!x}\,dx$ holds form (2.5)
The study of the integrals $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n}\!x}\,dx$ and $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n+1}\!x}\,dx$ relays solely on contour integration techniques. We made the remark that the Laurent expansion approach is very fruitful to find the residues here. To find these residues, we note that $\cosh^{n}\!z$ has a pole of $n$-th order at $z_{l}:=\dfrac{(2l+1)i\pi}{2}$ where $l\in\mathbb{Z}$. We find the Laurent expansion of $\dfrac{1}{\cosh^{n}\!z}{}$ around $z_{l}$ by letting first of all $z=z_{l}+w$. It comes
@@
In this paragraph, we mean by $\widetilde{\beta\zeta}(n)$ either the number $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ or $\dfrac{\beta(2n)}{\pi^{2n-1}}$ with $n\in\mathbb{N}^{*}$
@@
  • On putting the focus on the integral $\int_{0}^{1}\frac{x^{2p-1}}{\sqrt{1-x^{2}}\,\operatorname{arctanh}_{,}x}\,dx$, proposition 4.1 ensures the existence of some rational coefficients $r_{p,n}$ such that $\dfrac{\beta(2n)}{\pi^{2n-1}}=\displaystyle\sum_{p=1}^{n}r_{p,n}\displaystyle\int_{0}^{1}\frac{x^{2p-1}}{\sqrt{1-x^{2}}\,\operatorname{arctanh}_{,}x}\,dx$; thus, the existence of polynomials $\Xi_{n}\in\mathbb{Q}[x]$ such that $\dfrac{\beta(2n)}{\pi^{2n-1}}=\displaystyle\int_{0}^{1}\frac{x\,\Xi_{n}(x)}{\sqrt{1-x^{2}}\,\text{arctanh }x}\,dx$.
@@
  • On putting the focus on the integral $\int_{0}^{1}\frac{x^{2p-1}}{\operatorname{arctanh}_{,}x}\,dx$, proposition 4.1 ensures the existence of some rational coefficients $s_{p,n}$ such that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\displaystyle\sum_{p=1}^{n}s_{p,n}\displaystyle\int_{0}^{1}\frac{x^{2p-1}}{\operatorname{arctanh}_{,}x}\,dx$; thus, the existence of polynomials $\Lambda_{n}\in\mathbb{Q}[x]$ such that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\displaystyle\int_{0}^{1}\frac{x\,\Lambda_{n}(x)}{\text{arctanh }x}\,dx$.
@@
where $\psi$ is a normal common function. The rationality of the ratios $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ would be surprising. And the same line of reasoning holds for $\dfrac{\beta(2n)}{\pi^{2n}}$

### v2

@@ Introduction
Given the futility of the extraordinarily numerous attempts to compute these numbers, the solution of the designated problem is generally regarded as impossible; but so far a strong proof of this impossibility is lacking. Only heuristic methods and deep computations by brute force support this impossibility to this day. One may solve the problem by proving that the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ and $\dfrac{\beta(2n)}{\pi^{2n-1}}$ are not multiples of $\pi$. The difficulty of this new reformulation partly stems from the lack of sufficiently many closed-form expressions or identities involving these values that might serve as a starting point for an irrationality or transcendence proof. This paper is devoted to some integral representations of the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ and $\dfrac{\beta(2n)}{\pi^{2n-1}}$ through interesting logarithmic, trigonometric and hyperbolic kernels, along with several conjectures that support their non representability as rational multiples of $\pi$.
@@
The proof of this statement is fully detailed in [3] and is omitted here in the interest of brevity. We merely note the existence of integrals $I_{n}$ satisfying
@@
The formal existence of integrals of the type (2.4) or (2.5) can already be traced back to the work of Tobias Kyrion [11], where the author investigates, among others, the following integrals:
$\int_{0}^{+\infty}\frac{\tanh x}{x}\,\operatorname{sech}^{2n+1}\!x\,dx,\qquad\int_{0}^{+\infty}\frac{\tanh x}{x}\,\operatorname{sech}^{2n}\!x\,dx,\qquad\int_{0}^{+\infty}\frac{\tanh^{n}x}{x^{n}}\,dx.$
These three families of integrals were later rediscovered in [3] within a broader analytic framework, and are adopted here in a different formulation that is more suitable for the evaluation method and for discussing certain linear independence conjectures. Moreover, several additional integrals also satisfy either (2.4) or (2.5). The collection of all such integrals, together with the remarkable structural patterns they exhibit, constitutes the main contribution of this article. A representative selection is summarized in the following table, where $n$ denotes a strictly positive integer and $C_{p,n}$ are rational coefficients with $p\in\llbracket 1,n\rrbracket$.
@@
The reader is referred to [3] for illustrative examples of these integrals and for detailed derivations. We remark in passing that the integral $\int_{0}^{1}\frac{\operatorname{Im}\!\left[\operatorname{Li}_{-2n}(ix)\right]}{x}\,\ln\ln\tfrac{1}{x}\,dx$ generalizes the integral
@@
that has been firstly evaluated by Adamchik [5] and rediscovered by Blagouchine [4]. Additionally, the integral $\int_{0}^{1}\frac{\operatorname{Li}_{-2n-1}(-x^{2})}{x}\,\ln\ln\tfrac{1}{x}\,dx$ exhibits a general form of $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ by also extending the Blagouchine’s integral $\int_{0}^{1}\dfrac{x(x^{4}-4x^{2}+1)\ln\ln\frac{1}{x}}{(1+x^{2})^{4}}dx=\frac{7\,\zeta(3)}{8\,\pi^{2}}$ to all higher $n$. Strictly speaking, the endeavour to generalize the integral representations of Adamchik and Blagouchine did not originate with the present work. Indeed, Tobias Kyrion, in his article [11], establishes the following formulae, where $N$ denotes a strictly positive integer:
$\beta(2N)$ $=(-1)^{N}\,\frac{\pi^{\,2N-1}}{2^{\,2N-1}(2N-1)!}\int_{1}^{\infty}\sum_{k=0}^{N-1}\left(\sum_{j=0}^{k}\binom{2k+1}{k-j}(-1)^{\,j+1}(2j+1)^{\,2N-1}\right)$
$\quad\times\frac{\bigl((2k+1)x^{4}-2(2k+3)x^{2}+2k+1\bigr)\,x^{2k}}{(x^{2}+1)^{2k+3}}\log(\log(x))\,dx.$
$\zeta(2N+1)$ $=(-1)^{N}\,\frac{2\pi^{\,2N}}{(2^{\,2N+1}-1)\,(2N)!}\int_{1}^{\infty}\sum_{k=1}^{N}\left(\sum_{j=1}^{k}\binom{2k}{k-j}(-1)^{j}(2j)^{\,2N}\right)$
$\quad\times\frac{\bigl(2kx^{4}-2(2k+2)x^{2}+2k\bigr)\,x^{2k-1}}{(x^{2}+1)^{2k+2}}\log(\log(x))\,dx.$
Although these formulae are elegant and reveal the underlying structure of the integral, they involve multiple nested summations together with binomial coefficients, which makes them rather cumbersome to handle in irrationality proofs. The general integrals listed in the table above, by contrast, yield new representations that avoid such complications entirely, as the only special function that appears is the well-known Jonquière’s polylogarithm. Closed-form expressions for these new integrals are provided in the following propositions.
@@
with rational coefficients. Derivations steps detailed in [3] and omitted here for the sake of brevity. The following lines only give instructions on how to prove that
@@
holds form (2.4) and $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n}\!x}\,dx$ holds form (2.5)
The study of the integrals $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n}\!x}\,dx$ and $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n+1}\!x}\,dx$ relays solely on contour integration techniques by bearing in mind first of all the evenness of the integrands which imply that the integration bounds can be extended to $\infty$$\infty$[. We made the remark that the Laurent expansion approach is very fruitful to find the residues here. To find these residues, we note that $\cosh^{n}\!z$ has a pole of $n$-th order at $z_{l}:=\dfrac{(2l+1)i\pi}{2}$ where $l\in\mathbb{Z}$. We find the Laurent expansion of $\dfrac{1}{\cosh^{n}\!z}{}$ around $z_{l}$ by letting first of all $z=z_{l}+w$. It comes
@@
In virtue of Proposition (2.6), the irrationality of $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ is equivalent to this conjecture :
The set
$\left\{\pi,\ \int_{0}^{1}\frac{\operatorname{Li}_{-2n-1}(-x^{2})}{x}\,\ln\ln\frac{1}{x}\,dx\ \right\}$
is linearly independent over $\mathbb{Z}$ with $n\geq 1$
In virtue of Proposition (2.7), the irrationality of $\dfrac{\beta(2n)}{\pi^{2n}}$ is equivalent to this conjecture :
The set
$\left\{\pi,\ \int_{0}^{1}\frac{\operatorname{Im}\!\bigl[\operatorname{Li}_{-2n}(i\,x)\bigr]}{x}\,\ln\ln\frac{1}{x}\,dx\ \right\}$
is linearly independent over $\mathbb{Z}$ with $n\geq 1$
In further lines of this paragraph, we mean by $\widetilde{\beta\zeta}(n)$ either the number $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ or $\dfrac{\beta(2n)}{\pi^{2n-1}}$ with $n\in\mathbb{N}^{*}$
@@
  • On putting the focus on the integral $\int_{0}^{1}\frac{x^{2p-1}}{\sqrt{1-x^{2}}\,\operatorname{arctanh}\,x}\,dx$, proposition 4.1 ensures the existence of some rational coefficients $r_{p,n}$ such that $\dfrac{\beta(2n)}{\pi^{2n-1}}=\displaystyle\sum_{p=1}^{n}r_{p,n}\displaystyle\int_{0}^{1}\frac{x^{2p-1}}{\sqrt{1-x^{2}}\,\operatorname{arctanh}\,x}\,dx$; thus, the existence of polynomials $\Xi_{n}\in\mathbb{Q}[x]$ such that $\dfrac{\beta(2n)}{\pi^{2n-1}}=\displaystyle\int_{0}^{1}\frac{x\,\Xi_{n}(x)}{\sqrt{1-x^{2}}\,\text{arctanh }x}\,dx$.
@@
  • On putting the focus on the integral $\int_{0}^{1}\frac{x^{2p-1}}{\operatorname{arctanh}\,x}\,dx$, proposition 4.1 ensures the existence of some rational coefficients $s_{p,n}$ such that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\displaystyle\sum_{p=1}^{n}s_{p,n}\displaystyle\int_{0}^{1}\frac{x^{2p-1}}{\operatorname{arctanh}\,x}\,dx$; thus, the existence of polynomials $\Lambda_{n}\in\mathbb{Q}[x]$ such that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\displaystyle\int_{0}^{1}\frac{x\,\Lambda_{n}(x)}{\text{arctanh }x}\,dx$.
@@
where $\psi$ is a normal common function. The rationality of the ratios $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ and $\dfrac{\beta(2n)}{\pi^{2n}}$ would be very surprising.